# Likelihood-Ratio Regularized Quantile Regression: Adapting Conformal Prediction to High-Dimensional Covariate Shifts

### v1

@@ 4.1 Infinite sample setting
We first consider the infinite sample or “population” setting, characterizing the solutions of the LR-QR problem from (2) in an idealized scenario where the exact values of the expectations $\mathbb{E}_{1},\mathbb{E}_{2}$ can be calculated. In this case, we will show that if the function class $\mathcal{H}$ is linear and contains the true likelihood ratio $r$, then the optimizer achieves valid coverage in the test domain. Let $r_{\mathcal{H}}$ be the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. The key step is the result below, which characterizes coverage weighted by $r_{\mathcal{H}}$.
@@
Let $\mathcal{H}$ be a linear function class consisting of square-integrable functions with respect to $\mathbb{P}_{1,X}$. Then under regularity conditions specified in Appendix E (the conditions of Lemma L.3), if $(h^{*},\beta^{*})$ is a minimizer of the objective in Equation LR-QR, then we have
$\mathbb{E}_{1}[r_{\mathcal{H}}(X)\mathbf{1}[S\leqslant h^{*}(X)]]\geqslant 1-\alpha.$
The proof is given in Appendix I. As a consequence of Proposition 4.1, if $\mathcal{H}$ contains the true likelihood ratio $r$, so that $r_{\mathcal{H}}=r$, then in the infinite sample setting, the LR-QR threshold function $h^{*}$ attains valid coverage at test-time:
$\mathbb{E}_{1}[r(X)\mathbf{1}[S\leqslant h^{*}(X)]]=\mathbb{P}_{2}[S\leqslant h^{*}(X)]\geqslant 1-\alpha.$
However, in practice, we can only optimize over finite-dimensional function classes, and as a result we must control the effect of mis-specifying $\mathcal{H}$. If $r$ is not in $\mathcal{H}$, we can derive a lower bound on the coverage as follows. First, write
@@
By Proposition 4.1, the first term on the right-hand side is at least $1-\alpha$. Since the random variable $\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]$ is $\{0,1\}$-valued, the Cauchy-Schwarz inequality implies that the second term on the right-hand side is at least $-\mathbb{E}_{1}[(r-r_{\mathcal{H}})^{2}]^{1/2}$. Thus we have the lower bound
$\mathbb{P}_{2}[Y\in C^{*}(X)]$ $=\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]$
$\geqslant(1-\alpha)-\mathbb{E}_{1}[(r-r_{\mathcal{H}})^{2}]^{1/2}.$
Geometrically, this coverage gap is the result of restricting to $\mathcal{H}$; in fact, $\mathbb{E}_{1}[(r-r_{\mathcal{H}})^{2}]^{1/2}$ is the distance from $r$ to $\mathcal{H}$. This error decreases if $\mathcal{H}$ is made larger, but in the finite sample setting, this comes at the risk of overfitting.
@@ 4.2 Finite sample setting
From the analysis of the infinite sample regime, it is clear that if the function class $\mathcal{H}$ is larger, the test-time coverage of the population level LR-QR threshold function $h^{*}$ moves closer to the nominal value. However, in the finite sample setting, optimizing over a larger function class also presents the risk of overfitting. In a finite sample setting, by tuning the regularization parameter $\lambda$, we are trading off the estimation error incurred for the first term of Equation 2, namely $(\hat{\mathbb{E}}_{1}-\mathbb{E}_{1})[\ell_{\alpha}(h(X),S)]$, and the error incurred for the second and third terms of Equation 2, namely $\lambda(\hat{\mathbb{E}}_{3}-\mathbb{E}_{3})[\beta^{2}h^{2}]+\lambda(\hat{\mathbb{E}}_{2}-\mathbb{E}_{2})[-2\beta h]$. Heuristically, for a fixed $h$, the former should be proportional to $1/\sqrt{n_{1}}$, where $n_{1}$ is the number of labeled calibration datapoints, and the latter should be proportional to $\lambda(1/\sqrt{n_{3}}+1/\sqrt{n_{2}})$. Thus, if we pick the regularization to make these two errors of equal order, it will be proportional to $\sqrt{(n_{2}+n_{3})/n_{1}}$.
@@
Put differently, in order to ensure that the Empirical LR-QR threshold $\hat{h}$ from Equation Empirical-LR-QR has valid test coverage, one must choose the regularization $\lambda$ based on the relative amount of labeled and unlabeled data. The unlabeled datapoints carry information about the covariate shift $r$, because $r$ depends only on the distribution of the features. The labeled datapoints provide information about the conditional $(1-\alpha)$-quantile function $q_{1-\alpha}$, which depends only on the conditional distribution of $S|X$. When $\lambda$ is large, our optimization problem places more weight on approximating $r$ (the minimizer of $\mathbb{E}_{1}[(\beta h-r)^{2}]$ in $\beta h$), and if $\lambda$ is small, we instead aim to approximate $q_{1-\alpha}$ (the minimizer of $\ell_{\alpha}(h,S)$ in $h$). Therefore, if the number of unlabeled datapoints ($n_{2}+n_{3}$) is large compared to the number of labeled datapoints ($n_{1}$), our data contains much more information about the covariate shift $r$, and we should set $\lambda$ to be large. If instead $n_{1}$ is very large, the quantile function $q_{1-\alpha}$ can be well-approximated from the labeled calibration datapoints alone, and we set $\lambda$ to be close to zero. In the theoretical results, we make this intuition precise.
@@
In order to facilitate our theoretical analysis in the finite sample setting, we consider constrained versions of Equation LR-QR and Equation Empirical-LR-QR. Fix a collection $\Phi=(\phi_{1},\ldots,\phi_{d})^{\top}$ of $d$ basis functions, where $\phi_{i}:\mathcal{X}\to\mathbb{R}$ for $i\in[d]$. Let $\mathcal{I}=[\beta_{\textnormal{min}},\beta_{\textnormal{max}}]\subset\mathbb{R}$ be a compact interval with $\beta_{\textnormal{min}}>0$. Let $\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ be the $B$-ball centered at the origin in the linear function class spanned by $\{\phi_{1},\ldots,\phi_{d}\}$. We equip $\mathcal{H}_{B}$ with the norm $\|h\|=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$.
At the population level, consider the following constrained scaled LR-QR problem:
$(h^{*},\beta^{*})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}L(h,\beta).$ (4)
Also consider the following empirical constrained scaled LR-QR problem:
$(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}\hat{L}(h,\beta).$ (5)
@@
Under the regularity conditions specified in Appendix E, and for appropriate choices of the optimization hyperparameters [^2] Specifically, suppose that $\beta_{\textnormal{min}}\leqslant\beta_{\textnormal{lower}}$, $\beta_{\textnormal{max}}\geqslant\beta_{\textnormal{upper}}$, and $B\geqslant B_{\textnormal{upper}}$, where the positive scalars $\beta_{\textnormal{lower}}$, $\beta_{\textnormal{upper}}$, and $B_{\textnormal{upper}}$ are defined in Lemma L.4 in the Appendix, and depend on the data distribution and the choice of basis functions, but not on the data, the sample sizes, or the regularization parameter $\lambda$., for sufficiently large $n_{1},n_{2},n_{3}$, with probability at least $1-\delta$, any optimizer $(\hat{h},\hat{\beta})$ of the empirical constrained LR-QR objective from (5) has suboptimality gap $L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})$ with respect to the population risk (2) bounded by
$\mathcal{E}_{\textnormal{gen}}:=c\lambda\sqrt{\frac{1}{n_{2}}+\frac{1}{n_{3}}}+\frac{c^{\prime}}{\sqrt{n_{1}}}+\frac{c^{\prime\prime}}{\sqrt{\lambda}}\frac{1}{\sqrt{n_{1}}},$
@@
The proof is in Appendix J. The generalization error $\mathcal{E}_{\textnormal{gen}}$ is minimized for an optimal regularization on the order of
$\lambda^{*}\propto\frac{1}{n_{1}^{1/3}}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{-1/3},$ (6)
@@
which yields an optimized upper bound of order
$\mathcal{E}_{\textnormal{gen}}^{*}=O\left(\frac{1}{n_{1}^{1/3}}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/6}+\frac{1}{\sqrt{n_{1}}}\right).$
@@
As a corollary of Theorem 4.2, we have the following lower bound on the excess marginal coverage of our ERM threshold $\hat{h}$ in the covariate shifted domain. Let $r_{\mathcal{H}_{B}}$ denote the projection of $r$ onto the closed convex set $\mathcal{H}_{B}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$.
Theorem 4.3 (Lower bound on marginal coverage of ERM threshold under covariate shift).
@@
Under the same conditions as Theorem 4.2, consider the LR-QR optimizers $\hat{h}$ and $\hat{\beta}$ from (5). Given any $\delta>0$, for sufficiently large $n_{1},n_{2},n_{3}$, we have with probability at least $1-\delta$ that
@@
$\mathbb{P}_{2}\left[Y\in\hat{C}(X)\right]$ $\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|],$
where
$\mathcal{E}_{\textnormal{cov}}:=A\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\lambda+A^{\prime}\frac{\lambda^{-1/4}}{n_{1}^{1/4}}+\Theta\left(\frac{\lambda^{1/2}}{n_{1}^{1/4}}\right),$
@@
and $A,A^{\prime}$ are positive scalars that do not depend on $\lambda$.
@@
The proof is in Appendix K. This result states that our LR-QR method has nearly valid coverage at level $1-\alpha$ under covariate shift, up to small error terms that we can control. The quantity $\mathcal{E}_{\textnormal{cov}}$ vanishes as we collect more data. The term $\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|]$ captures the level of mis-specification by not including the true likelihood ratio function $r$ in our function class $\mathcal{H}_{B}$. This can be decreased by making the function class $\mathcal{H}_{B}$ larger. Of course, this will also increase the size of the terms $A,A^{\prime}$ in our coverage error, but in our theory we show that the dependence is mild. Indeed, the terms depend only on a few geometric properties of $\mathcal{H}_{B}$ such as the eigenvalues of the sample covariance matrix of the basis $\Phi(X)$ under the source distribution, and a quantitative measure of linear dependence of the features; but not explicitly on the dimension of the basis.
@@
We highlight the term $2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]$, which is an error term relating the projected likelihood ratio $r_{\mathcal{H}_{B}}$ to the LR-QR solution $\hat{\beta}\hat{h}$. Crucially, this term is a non-negative quantity multiplied by $\lambda$, and so for appropriate $\lambda$ it may counteract in part the coverage error loss. Consistent with the above observations, we find empirically that choosing small nonzero regularization parameters improves coverage. Moreover, we find that choosing the regularization parameter to be on the order of the optimal value for $\mathcal{E}_{\textnormal{cov}}$ is suitable choice across a range of experiments.
@@
Our proofs are quite involved and require a number of delicate arguments. Crucially, they draw on a novel analysis of coverage via stability bounds from learning theory. Existing stability results cannot directly be applied to our optimization problem, due to our use of a data-dependent regularizer. For instance, in classical settings, the optimal regularization tends to zero as the sample size goes to infinity, but this is not the case for our optimization problem. To overcome this challenge, we combine stability bounds Shalev-Shwartz et al., (2010); Shalev-Shwartz and Ben-David, (2014) with a novel conditioning argument, and we show that the values of $L$ at the minimizers of $\hat{L}$ and $L$ are close by introducing intermediate losses that sequentially swap out empirical expectations $\hat{\mathbb{E}}_{1},\hat{\mathbb{E}}_{2},\hat{\mathbb{E}}_{3}$ with their population counterparts. We then leverage the smoothness of $L$, to derive that the gradient of $L$ at $(\hat{\beta},\hat{h})$ is small. Expanding on our discussion from Section 3.2, we show that a small gradient implies the desired small coverage gap. Our methods may be of independent interest for deriving generalization bounds of empirical risk minimizers with data-dependent regularizers.

### v2

@@ 4.1 Infinite sample setting
We first consider the infinite sample or “population” setting, characterizing the solutions of the LR-QR problem from (LR-QR) in an idealized scenario where the exact values of the expectations $\mathbb{E}_{1},\mathbb{E}_{2}$ can be calculated. In this case, we will show that if the hypothesis class $\mathcal{H}$ is linear and contains the true likelihood ratio $r$, then the optimizer achieves valid coverage in the test domain. Let $r_{\mathcal{H}}$ be the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. The key step is the result below, which characterizes coverage weighted by $r_{\mathcal{H}}$.
@@
Let $\mathcal{H}$ be a linear hypothesis class consisting of square-integrable functions with respect to $\mathbb{P}_{1,X}$. Then under regularity conditions specified in Appendix F (the conditions of Lemma M.3), if $(h^{*},\beta^{*})=(h^{*}_{\lambda},\beta^{*}_{\lambda})$ is a minimizer of the objective in Equation LR-QR with regularization strength $\lambda>0$, then we have $\mathbb{E}_{1}[r_{\mathcal{H}}(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]\geqslant 1-\alpha.$
The proof is given in Appendix J. As a consequence of Proposition 4.1, if $\mathcal{H}$ contains the true likelihood ratio $r$, so that $r_{\mathcal{H}}=r$, then in the infinite sample setting, the LR-QR threshold function $h^{*}$ attains valid coverage at test-time:
$\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]=\mathbb{P}_{2}[S(X,Y)\leqslant h^{*}(X)]\geqslant 1-\alpha.$
However, in practice, we can only optimize over finite-dimensional hypothesis classes, and as a result we must control the effect of mis-specifying $\mathcal{H}$. If $r$ is not in $\mathcal{H}$, we can derive a lower bound on the coverage as follows. First, write
@@
By Proposition 4.1, the first term on the right-hand side is at least $1-\alpha$. Since the random variable $\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]$ is $\{0,1\}$-valued, the Cauchy-Schwarz inequality implies that the second term on the right-hand side is at least $-\mathbb{E}_{1}[(r(X)-r_{\mathcal{H}}(X))^{2}]^{1/2}$. We set our threshold function $q$ to equal $h^{*}$, so that our conformal prediction sets equal $C^{*}(x)=\{y\in\mathcal{Y}:S(x,y)\leqslant h^{*}(x)\}$ for all $x\in\mathcal{X}$. Thus, we have the lower bound
$\mathbb{P}_{2}[Y\in C^{*}(X)]$ $=\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]\geqslant(1-\alpha)-\mathbb{E}_{1}[(r(X)-r_{\mathcal{H}}(X))^{2}]^{1/2}.$
Geometrically, this coverage gap is the result of restricting to $\mathcal{H}$; in fact, $\mathbb{E}_{1}[(r(X)-r_{\mathcal{H}}(X))^{2}]^{1/2}$ is the distance from $r$ to $\mathcal{H}$. This error decreases if $\mathcal{H}$ is made larger, but in the finite sample setting, this comes at the risk of overfitting.
@@ 4.2 Finite sample setting
From the analysis of the infinite sample regime, it is clear that if the hypothesis class $\mathcal{H}$ is made larger, the test-time coverage of the population level LR-QR threshold function $h^{*}$ moves closer to the nominal value. However, in the finite sample setting, optimizing over a larger hypothesis class also presents the risk of overfitting. By tuning the regularization parameter $\lambda$, we are trading off the estimation error incurred for the first term of Equation LR-QR, namely $(\hat{\mathbb{E}}_{1}-\mathbb{E}_{1})[\ell_{\alpha}(h(X),S(X,Y))]$, and the error incurred for the second and third terms of Equation LR-QR, namely $\lambda(\hat{\mathbb{E}}_{3}-\mathbb{E}_{3})[\beta^{2}h(X)^{2}]+\lambda(\hat{\mathbb{E}}_{2}-\mathbb{E}_{2})[-2\beta h(X)]$. Heuristically, for a fixed $h$, the former should be proportional to $1/\sqrt{n_{1}}$, and the latter should be proportional to $\lambda(1/\sqrt{n_{3}}+1/\sqrt{n_{2}})$. Thus, if we pick $\lambda$ to make these two errors of equal order, it will be proportional to $\sqrt{(n_{2}+n_{3})/n_{1}}$.
@@
Put differently, in order to ensure that the Empirical LR-QR threshold $\hat{h}$ from Equation Empirical-LR-QR has valid test coverage, one must choose the regularization $\lambda$ based on the relative amount of labeled and unlabeled data. The unlabeled datapoints carry information about the covariate shift $r$, because $r$ depends only on the distribution of the features. The labeled datapoints provide information about the conditional $(1-\alpha)$-quantile function $q_{1-\alpha}$, which depends only on the conditional distribution of $S|X$. When $\lambda$ is large, our optimization problem places more weight on approximating $r$ (the minimizer of $\mathbb{E}_{1}[(\beta h(X)-r(X))^{2}]$ in $\beta h$), and if $\lambda$ is small, we instead aim to approximate $q_{1-\alpha}$ (the minimizer of $\mathbb{E}_{1}[\ell_{\alpha}(h(X),S(X,Y))]$ in $h$). Therefore, if the number of unlabeled datapoints ($n_{2}+n_{3}$) is large compared to the number of labeled datapoints ($n_{1}$), our data contains much more information about the covariate shift $r$, and we should set $\lambda$ to be large. If instead $n_{1}$ is very large, the quantile function $q_{1-\alpha}$ can be well-approximated from the labeled calibration datapoints, and we set $\lambda$ to be close to zero. In the theoretical results, we make this intuition precise.
@@
In order to facilitate our theoretical analysis in the finite sample setting, we consider constrained versions of Equation LR-QR and Equation Empirical-LR-QR. Fix a collection $\Phi=(\phi_{1},\ldots,\phi_{d})^{\top}$ of $d$ basis functions, where $\phi_{i}:\mathcal{X}\to\mathbb{R}$ for $i\in[d]$. Let $\mathcal{I}=[\beta_{\textnormal{min}},\beta_{\textnormal{max}}]\subset\mathbb{R}$ be an interval with $\beta_{\textnormal{min}}>0$. Let $\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ be the $B$-ball centered at the origin in the linear hypothesis class spanned by $\{\phi_{1},\ldots,\phi_{d}\}$. We equip $\mathcal{H}_{B}$ with the norm $\|h\|=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$.
At the population level, consider the following constrained LR-QR problem: $(h^{*},\beta^{*})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}L_{\lambda}(h,\beta)$. Also consider the following empirical constrained LR-QR problem [^2] For brevity, this notation overloads the definition of $(\hat{h},\hat{\beta})$ from (Empirical-LR-QR). From now on, $(\hat{h},\hat{\beta})$ will refer to the definition from (3), and the one from (Empirical-LR-QR) will not be used again.:
$(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}\hat{L}_{\lambda}(h,\beta).\vspace{-0.5em}$ (3)
@@
Under the regularity conditions specified in Appendix F, and for appropriate choices of the optimization hyperparameters [^3] Specifically, suppose that $\beta_{\textnormal{min}}\leqslant\beta_{\textnormal{lower}}$, $\beta_{\textnormal{max}}\geqslant\beta_{\textnormal{upper}}$, and $B\geqslant B_{\textnormal{upper}}$, where the positive scalars $\beta_{\textnormal{lower}}$, $\beta_{\textnormal{upper}}$, and $B_{\textnormal{upper}}$ are defined in Lemma M.4 in the Appendix, and depend on the data distribution and the choice of basis functions, but not on the data, the sample sizes, or the regularization parameter $\lambda$., for sufficiently large $n_{1},n_{2},n_{3}$, with probability at least $1-\delta$, any optimizer $(\hat{h},\hat{\beta})$ of the empirical constrained LR-QR objective from (3) with regularization strength $\lambda>0$ has suboptimality gap $L_{\lambda}(\hat{h},\hat{\beta})-L_{\lambda}(h^{*},\beta^{*})$ with respect to the population risk (LR-QR) bounded by
$\mathcal{E}_{\textnormal{gen}}:=c\lambda\sqrt{{1}/{n_{2}}+{1}/{n_{3}}}+{c^{\prime}}/{\sqrt{n_{1}}}+{c^{\prime\prime}}/{\sqrt{\lambda n_{1}}},$
@@
The proof is in Appendix K. The generalization error $\mathcal{E}_{\textnormal{gen}}$ is minimized for an optimal regularization on the order of
$\lambda^{*}\propto{n_{1}^{-1/3}}\left({1}/{n_{2}}+{1}/{n_{3}}\right)^{-1/3},$ (4)
@@
which yields an optimized upper bound of order $\mathcal{E}_{\textnormal{gen}}^{*}=O\left({n_{1}^{-1/3}}\left({1}/{n_{2}}+{1}/{n_{3}}\right)^{1/6}+1/\sqrt{n_{1}}\right)$.
@@
As a corollary of Theorem 4.2, we have the following lower bound on the excess marginal coverage of our ERM threshold $\hat{h}$ in the covariate shifted domain. Let $r_{B}$ denote the projection of $r$ onto the closed convex set $\mathcal{H}_{B}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$.
Theorem 4.3 (Main result: Coverage under covariate shift).
@@
Under the same conditions as Theorem 4.2, consider the LR-QR optimizers $\hat{h}$ and $\hat{\beta}$ from (3) with regularization strength $\lambda>0$. Given any $\delta>0$, for sufficiently large $n_{1},n_{2},n_{3}$, we have with probability at least $1-\delta$ that [^4] The probability $\mathbb{P}_{2}\left[Y\in\hat{C}(X)\right]$ is over $(X,Y)\sim\mathbb{P}_{2}$, conditional on $\hat{C}$.
@@
$\mathbb{P}_{2}\left[Y\in\hat{C}(X)\right]$ $\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{B}(X)-\hat{\beta}\hat{h}(X))^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)\mathbb{E}_{1}[|r(X)-r_{B}(X)|],$
@@
where $\mathcal{E}_{\textnormal{cov}}:=A\left({1}/{n_{2}}+{1}/{n_{3}}\right)^{1/4}\lambda+A^{\prime}(\lambda n_{1})^{-1/4}+\lambda^{1/2}/n_{1}^{1/4},$ and $A,A^{\prime}$ are positive scalars that do not depend on $\lambda$.
@@
The proof is in Appendix L. This result states that our LR-QR method has nearly valid coverage at level $1-\alpha$ under covariate shift, up to small error terms that we can control. The quantity $\mathcal{E}_{\textnormal{cov}}$ vanishes as we collect more data. The term $\mathbb{E}_{1}[|r(X)-r_{B}(X)|]$ captures the level of mis-specification by not including the true likelihood ratio function $r$ in our hypothesis class $\mathcal{H}_{B}$. This can be decreased by making the hypothesis class $\mathcal{H}_{B}$ larger. Of course, this will also increase the size of the terms $A,A^{\prime}$ in our coverage error, but in our theory we show that the dependence is mild. Indeed, the terms depend only on a few geometric properties of $\mathcal{H}_{B}$ such as the eigenvalues of the sample covariance matrix of the basis $\Phi(X)$ under the source distribution, and a quantitative measure of linear dependence of the features; but not explicitly on the dimension of the basis.
@@
We highlight the term $2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{B}(X)-\hat{\beta}\hat{h}(X))^{2}]$, which is an error term relating the projected likelihood ratio $r_{B}$ to the LR-QR solution $\hat{\beta}\hat{h}$. Crucially, this term is a non-negative quantity multiplied by $\lambda$, and so for appropriate $\lambda$ it may counteract in part the coverage error loss. Consistent with the above observations, we find empirically that choosing small nonzero regularization parameters improves coverage. Moreover, we find that choosing the regularization parameter to be on the order of the optimal value for $\mathcal{E}_{\textnormal{cov}}$ is suitable choice across a range of experiments.
@@
Our proofs are quite involved and require a number of delicate arguments. Crucially, they draw on a novel analysis of coverage via stability bounds from learning theory. Existing stability results cannot directly be applied, due to our use of a data-dependent regularizer. For instance, in classical settings, the optimal regularization tends to zero as the sample size goes to infinity, but this is not the case here. To overcome this challenge, we combine stability bounds Shalev-Shwartz et al., (2010); Shalev-Shwartz and Ben-David, (2014) with a novel conditioning argument, and we show that the values of $L$ at the minimizers of $\hat{L}$ and $L$ are close by introducing intermediate losses that sequentially swap out empirical expectations $\hat{\mathbb{E}}_{1},\hat{\mathbb{E}}_{2},\hat{\mathbb{E}}_{3}$ with their population counterparts. We then leverage the smoothness of $L$, to derive that the gradient of $L$ at $(\hat{\beta},\hat{h})$ is small. Finally, we show that a small gradient implies the desired small coverage gap.
As an immediate corollary of Theorem 4.3, we have the following result, which states that the LR-QR algorithm can be used to construct prediction sets with group-conditional coverage for a finite set of potentially overlapping groups.
Corollary 4.4 (Group-conditional coverage).
Let $G_{1},\ldots,G_{d}\subseteq\mathcal{X}$ be a finite collection of distinct subsets of $\mathcal{X}$ such that $\mathbb{P}_{1}\left[G_{i}\right]>0$ for all $i\in[d]$ and $\mathbb{P}_{1}\left[G_{i}\triangle G_{j}\right]>0$ for all $i,j\in[d]$ with $i\neq j$, where $\triangle$ denotes symmetric difference. For $i\in[d]$, let $\phi_{i}:\mathcal{X}\to\mathbb{R}$ be given by $\phi_{i}(x)=\mathbf{1}[x\in G_{i}]$, and consider the basis $\Phi=(\phi_{1},\ldots,\phi_{d})^{\top}$. Under the same conditions as Theorem 4.2, consider the LR-QR optimizers $\hat{h}$ and $\hat{\beta}$ from (3) with basis given by $\Phi$ and regularization strength $\lambda>0$. Given any $\delta>0$, for sufficiently large $n_{1},n_{2},n_{3}$, we have with probability at least $1-\delta$ that
$\mathbb{P}_{1}\left[Y\in\hat{C}(X)|X\in G_{i}\right]$ $\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(c_{i}\phi_{i}(X)-\hat{\beta}\hat{h}(X))^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)(1-c_{i}\mathbb{P}_{1}\left[G_{i}\right])$
for each $i\in[d]$, where $c_{i}=\min\{1/\mathbb{P}_{1}\left[G_{i}\right],B\}$ and $\mathcal{E}_{\textnormal{cov}}$ is defined as in Theorem 4.3.